# Corks, exotic 4-manifolds and genus functions

### v3

@@ 7. Genus functions and quasi-invertible cobordisms
In this section, we study behavior of genus functions of 4-manifolds with quasi-invertible cobordisms attached. We also show stabilities of algebraic inequivalences of genus functions under attachments of quasi-invertible cobordisms. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected unless otherwise stated, but we still assume surfaces are connected.
@@ 7.1. Behavior of genus functions
We first study behavior of genus functions under attachments of quasi-invertible cobordisms. Throughout this section, we use the following setting. Let $M,N$ be closed 3-manifolds, and let $P$ be a quasi-invertible cobordism from $M$ to $N$. Let $X$ be a connected 4-manifold having a boundary component orientedly diffeomorphic to $M$, and we denote a glued 4-manifold $X\cup_{M}P$ by $X^{\prime}$. Let $\iota:X\hookrightarrow X^{\prime}$ be the inclusion.
@@
(1) We first consider the case where the attached handle is a 3-handle $D^{3}\times D^{1}$. Let $j:Z\hookrightarrow Z^{\prime}$ denote the inclusion. We note that the kernel of $j_{*}:H_{2}(Z)\to H_{2}(Z^{\prime})$ is the image of the homomorphism $H_{2}(S^{2}\times D^{1})\to H_{2}(Z)$ induced by the inclusion as seen from the Mayer–Vietoris exact sequence, where $S^{2}\times D^{1}$ denotes the attaching region of the 3-hanlde.
@@
In the case where the attached handle is a 4-handle, the claim follows by a much simpler argument.
@@
Now we prove Theorem 7.1.
@@
We attach a quasi-inverse cobordism $Q$ of $P$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism. Let $X^{\prime\prime}$ denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$. The cobordism $R$ becomes diffeomorphic to $I\times M$ after performing surgeries along finitely many embedded 4-dimensional 1-handlebodies. Thus, the 4-manifold $\widehat{X}^{\prime\prime}$ obtained from $X^{\prime\prime}$ by these surgeries is diffeomorphic to $X\cup_{\mathrm{id}_{M}}(I\times M)$. We denote by $\widehat{j}:X\hookrightarrow\widehat{X}^{\prime\prime}$ the inclusion. By shrinking $I\times M$ together with the collar neighborhood of $\partial X$, we can easily construct a diffeomorphism $\varphi:\widehat{X}^{\prime\prime}\to X$ such that $\varphi\circ\widehat{j}:X\to X$ is homotopic to $\mathrm{id}_{X}$. Since the diffeomorphisms $\varphi$ and $\mathrm{id}_{X}$ preserves the genus functions, it follows that $\widehat{j}:X\hookrightarrow\widehat{X}^{\prime\prime}$ also preserves the genus functions.
We remove the interior of the aforementioned 1-handlebodies in $R$ from $X^{\prime\prime}$ and denote the resulting 4-manifold by $X^{\prime\prime}_{c}$. Let $j:X\hookrightarrow X^{\prime\prime}$, $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$, $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$, $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ and $\widehat{h}:X^{\prime\prime}_{c}\hookrightarrow\widehat{X}^{\prime\prime}$ be the inclusions. Since $j_{c}$ and $\widehat{h}$ are inclusions, we see that the inequalities $g_{X}(\alpha)\geq g_{X^{\prime\prime}_{c}}({j_{c}}_{*}(\alpha))$ and $g_{X^{\prime\prime}_{c}}(\beta)\geq g_{\widehat{X}^{\prime\prime}}(\widehat{h}_{*}(\beta))$ hold for any $\alpha\in H_{2}(X)$ and $\beta\in H_{2}(X^{\prime\prime}_{c})$. As noted before, $\widehat{j}=\widehat{h}\circ j_{c}:X\hookrightarrow\widehat{X}^{\prime\prime}$ preserves the genus functions. The above inequalities thus imply that $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$ also preserves the genus functions. By Lemma 7.2, the inclusion $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ preserves the genus functions, and thus $j=h\circ j_{c}:X\hookrightarrow X^{\prime\prime}$ also preserves the genus functions. This fact implies that $\iota:X\hookrightarrow X^{\prime}$ preserves the genus functions due to the relation $j=j^{\prime}\circ\iota$. ∎
@@
(1) By Theorem 7.1 and Lemma 6.7, the genus functions are preserved under attachment of a 1-handle whose attaching region belongs to a connected boundary component and also under attachment of 2-handles along a strongly slice link with 0-framings.
(2) To prove a result similar to Theorem 7.1 for other invariants, we note that the following properties (i)–(iii) of genus functions are used in the above proof. (i) Every diffeomorphism preserves the genus functions. (ii) For any orientation preserving embedding, the homomorphism induced by the embedding does not strictly increase the values of their genus functions. (iii) Lemma 7.2.(1). Even when we do not use Lemma 7.2.(1), we can still prove Theorem 7.1 in the case where $P$ is an invertible cobordism as seen from the above proof.
@@
If ${\iota}_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism, then the genus function of ${X^{\prime}}$ is algebraically equivalent to that of ${X}$.
The corollary below provides sufficient conditions that $g_{X^{\prime}}$ is algebraically equivalent to $g_{X}$.
@@
Then the genus function of ${X^{\prime}}$ is algebraically equivalent to that of ${X}$.
@@
This is straightforward from Corollary 7.4, Lemma 6.11, and Proposition 6.15. ∎
In the case where $P$ is $H_{2}$-surjective, the genus function $g_{X^{\prime}}$ is completely determined by $g_{X}$ and $\iota_{*}$ due to Theorem 7.1 and Lemma 6.11. In the case where $P$ is not $H_{2}$-surjective, we give estimates on the values of $g_{X^{\prime}}$ for all classes. As in Subsection 6.2, let $B_{P}$ be the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$ induced by the inclusion, where $R$ is the quasi-product cobordism defined in the proof of Theorem 7.1. Also, let $\iota:X\hookrightarrow X^{\prime}$ and $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusions. Due to Lemma 6.11, we have $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(B_{P})$. We prove the following estimates for $g_{X^{\prime}}$.
Proposition 7.6.
The following inequality holds for any $\alpha\in H_{2}(X)$ and any $\beta\in B_{P}$.
@@
The right inequality is obvious, so we prove the left inequality. Let $X^{\prime\prime}$ be as in the proof of Theorem 7.1, and let $j:X\hookrightarrow X^{\prime\prime}$ and $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$ be the inclusions. By the definition of $B_{P}$, we see $j^{\prime}_{*}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))=j_{*}(\alpha)$, showing $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X^{\prime\prime}}(j_{*}(\alpha))$. Since $j_{*}$ preserves the genus functions due to Theorem 7.1, we obtain $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X}(\alpha)$. ∎
There are many examples of quasi-invertible cobordisms $P$ such that $g_{P}$ is the zero-map, and for such examples, $g_{X^{\prime}}$ is completely determined even when $P$ is not $H_{2}$-surjective due to Proposition 7.6.
Similarly to ordinary genus functions, torsion-free genus functions are preserved by the inclusions as easily seen from Proposition 7.6 and Lemma 6.11.
@@
The inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions.
@@ 7.2. Stabilities for algebraic inequivalences of genus functions
We next discuss stabilities for algebraic inequivalences of genus functions under attachments of quasi-invertible cobordisms.
In the rest of this section, we use the following setting. For $i=1,2$, let $M_{i},N_{i}$ be closed 3-manifolds, let $P_{i}$ be a quasi-invertible cobordism from $M_{i}$ to $N_{i}$, and let $Q_{i}$ be its quasi-inverse cobordism. Also, let $X_{i}$ be a connected 4-manifolds having a boundary component diffeomorphic to $M_{i}$, and we denote a glued 4-manifold $X_{i}\cup_{M_{i}}P_{i}$ by $X^{\prime}_{i}$. We attach $Q_{i}$ to $X_{i}^{\prime}$ so that the submanifold $R_{i}=P_{i}\cup_{N_{i}}Q_{i}$ is a quasi-product cobordism. Let $X^{\prime\prime}$ denote the resulting 4-manifold $X_{i}\cup_{M_{i}}P_{i}\cup_{N_{i}}Q_{i}$. Let $\iota:X_{i}\hookrightarrow X^{\prime}_{i}$ and $j:X_{i}\hookrightarrow X^{\prime\prime}_{i}$ be the inclusions. Furthermore, we define the subgroup ${\iota_{P_{i}}}_{*}(B_{P_{i}})$ of $H_{2}(X_{i}^{\prime})$ as in Subsection 6.2.
The theorem below shows a stability for algebraic inequivalence of genus functions.
@@
$(1)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime}_{i})$ is an isomorphism. Then the genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is an isomorphism. Then the torsion-free genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
@@
For general quasi-invertible cobordisms, the stability of algebraic inequivalences still holds under additional assumptions. See Propositions 6.14 and 6.15 and Lemma 6.16 for sufficient conditions that the assumption below holds.
@@
Suppose the intersection form of each $X_{i}$ is non-degenerate. Suppose further that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups ${\iota_{P_{1}}}_{*}(B_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(B_{P_{2}})$ are torsion-free and that each ${j_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime\prime}_{i})$ is injective. If the genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ are algebraically equivalent, then those of ${X_{1}}$ and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the restrictions of the genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ respectively to ${\iota_{P_{1}}}_{*}(B_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(B_{P_{2}})$ are also algebraically equivalent.
@@
Due to Propositions 5.3 and 5.2, the claim follows from Lemma 6.11, Propositions 6.14 and 7.6 and, Theorem 7.1. ∎
For torsion-free genus functions, a similar stability holds under a much simpler assumption.
@@
Suppose the intersection form of each $X_{i}$ is non-degenerate. If the torsion-free genus functions of ${X^{\prime}_{1}}$ and $X^{\prime}_{2}$ are algebraically equivalent, then those of ${X_{1}}$ and $X_{2}$ are algebraically equivalent.
@@
Since the intersection form of each $X_{i}$ is non-degenerate, we see that $\iota_{*}:H_{"}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is injective. Due to Proposition 5.4, the claim follows from Lemmas 6.11 and 6.13, Proposition 7.6 and Corollary 7.7. ∎

### v5

@@ 7. Genus functions and quasi-invertible cobordisms
In this section, we study behavior of genus functions of 4-manifolds under attachment of quasi-invertible cobordisms. We also show stabilities of algebraic inequivalences of genus functions under the attachments. Throughout this section, we allow manifolds, submanifolds, and their boundaries to be disconnected unless otherwise stated, but we still assume surfaces are connected.
@@ 7.1. Behavior of genus functions
Throughout this subsection, we use the following setting. Let $M,N$ be closed 3-manifolds, and let $P$ be a quasi-invertible cobordism from $M$ to $N$. Let $X$ be a connected 4-manifold having a boundary component orientedly diffeomorphic to $M$, and we denote a glued 4-manifold $X\cup_{M}P$ by $X^{\prime}$. Let $\iota:X\hookrightarrow X^{\prime}$ be the inclusion.
@@
(1) We prove the claim in the case where the attached handle is a 3-handle $D^{3}\times D^{1}$. In the case of a 4-handle, the claim follows by a much simpler argument. Let $j:Z\hookrightarrow Z^{\prime}$ denote the inclusion. We note that the kernel of $j_{*}:H_{2}(Z)\to H_{2}(Z^{\prime})$ is the image of the homomorphism $H_{2}(S^{2}\times D^{1})\to H_{2}(Z)$ induced by the inclusion as seen from the Mayer–Vietoris exact sequence, where $S^{2}\times D^{1}$ denotes the attaching region of the 3-handle.
@@
In the rest of this subsection, we use the following setting to prove Theorem 7.1 and its corollaries. We attach a quasi-inverse cobordism $Q$ of $P$ to $X^{\prime}$ so that the submanifold $R=P\cup_{N}Q$ is a quasi-product cobordism, and we denote the resulting 4-manifold $X\cup_{M}P\cup_{N}Q$ by $X^{\prime\prime}$.
@@
The cobordism $R$ becomes diffeomorphic to $I\times M$ possibly after performing surgeries along finitely many embedded 4-dimensional 1-handlebodies. Thus, the 4-manifold $\widehat{X}^{\prime\prime}$ obtained from $X^{\prime\prime}$ by these surgeries is diffeomorphic to $X\cup_{\mathrm{id}_{M}}(I\times M)$. We denote by $\widehat{j}:X\hookrightarrow\widehat{X}^{\prime\prime}$ the inclusion. By shrinking $I\times M$ together with the collar neighborhood of $\partial X$, we can easily construct a diffeomorphism $\varphi:\widehat{X}^{\prime\prime}\to X$ such that $\varphi\circ\widehat{j}:X\to X$ is homotopic to $\mathrm{id}_{X}$. Since the diffeomorphisms $\varphi$ and $\mathrm{id}_{X}$ preserves the genus functions, it follows that $\widehat{j}:X\hookrightarrow\widehat{X}^{\prime\prime}$ also preserves the genus functions.
We remove the interior of the aforementioned 1-handlebodies in $R$ from $X^{\prime\prime}$ and denote the resulting 4-manifold by $X^{\prime\prime}_{c}$. Note that $X^{\prime\prime}=\widehat{X}^{\prime\prime}=X^{\prime\prime}_{c}$ in the case where $R$ is a product cobordism. Let $\hookrightarrow X^{\prime\prime}, j:XX′′j^{\prime}$$^{\prime}\hookrightarrow X^{\prime\prime}$, $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$, $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ and $\widehat{h}:X^{\prime\prime}_{c}\hookrightarrow\widehat{X}^{\prime\prime}$ be the inclusions. Since $j_{c}$ and $\widehat{h}$ are inclusions, and $\widehat{j}=\widehat{h}\circ j_{c}$ preserves the genus functions, every $\alpha\in H_{2}(X)$ satisfies
$g_{X}(\alpha)\geq g_{X^{\prime\prime}_{c}}({j_{c}}_{*}(\alpha))\geq g_{\widehat{X}^{\prime\prime}}(\widehat{h}_{*}({j_{c}}_{*}(\alpha)))=g_{X}(\alpha),$
showing that $j_{c}:X\hookrightarrow X^{\prime\prime}_{c}$ also preserves the genus functions. By Lemma 7.2, the inclusion $h:X^{\prime\prime}_{c}\hookrightarrow X^{\prime\prime}$ preserves the genus functions, and thus $j=h\circ j_{c}:X\hookrightarrow X^{\prime\prime}$ also preserves the genus functions. Due to this fact and the relation $j=j^{\prime}\circ\iota$, we see that every $\alpha\in H_{2}(X)$ satisfies
$g_{X}(\alpha)\geq g_{X^{\prime}}({\iota}_{*}(\alpha))\geq g_{{X}^{\prime\prime}}(j^{\prime}_{*}({\iota}_{*}(\alpha)))=g_{X}(\alpha),$
showing that $\iota:X\hookrightarrow X^{\prime}$ preserves the genus functions. ∎
@@
(1) By Theorem 7.1 and Lemma 6.7, the genus functions are preserved under attachment of a 1-handle whose attaching region belongs to a connected boundary component and also under attachment of 2-handles along a strongly slice link with 0-framings. Applying this result, we can show non-sliceness for various knots in $S^{3}$.
(2) To prove a result similar to Theorem 7.1 for other invariants in Section 11, we note that the properties of genus functions used in the above proof are the following (i)–(iii). (i) Every (orientation-preserving) diffeomorphism preserves the genus functions. (ii) For any (orientation-preserving) embedding, the homomorphism induced by the embedding does not increase the values of genus functions. (iii) The property stated in Lemma 7.2.(1). We note that, in the case where $P$ is an invertible cobordism, the above proof of Theorem 7.1 does not require Lemma 7.2.(1).
(3) There are many examples of cobordisms that do not have the property of Theorem 7.1. For instance, many cobordisms do not have this property after taking a connected sum with $S^{2}\times S^{2}$. We can also construct examples of such homology cobordisms as follows. Take a (contractible) cork $(C,f)$ admitting a 4-manifold $Z$ such that its cork twist along $(C,f)$ changes the algebraic equivalent class of the genus function of $Z$. Note that there are various examples of such corks (e.g. [6], [9], [10], [37]). Then, $\overline{C}-\textnormal{int}\,D^{4}$ is a homology cobordism from $\partial C$ to $S^{3}$, and the attachment of this cobordism to $\overline{Z-\textnormal{int}\,C}$ using either $f$ or $\mathrm{id}_{\partial C}$ does not preserve the genus function of $\overline{Z-\textnormal{int}\,C}$ due to the assumption on $Z$.
@@
If ${\iota}_{*}:H_{2}(X)\to H_{2}(X^{\prime})$ is an isomorphism, then the genus function of $X^{\prime}$} is algebraically equivalent to that of X{X}.
The corollaries below provide sufficient conditions that $g_{X^{\prime}}$ is algebraically equivalent to $g_{X}$.
Corollary 7.5.
If $P$ is an invertible homology cobordism from $M$ to $N$, then the genus function of $X^{\prime}$} is algebraically equivalent to that of X{X}.
Proof.
This is straightforward from Corollary 7.4. ∎
@@
Then the genus function of $X^{\prime}$} is algebraically equivalent to that of X{X}.
@@
This is straightforward from Corollary 7.4, Lemma 6.11, and Propositions 6.14 and 6.13. ∎
In the case where $P$ is $H_{2}$-surjective, the genus function $g_{X^{\prime}}$ is completely determined by $g_{X}$ and $\iota_{*}$ due to Theorem 7.1 and Lemma 6.11. In the case where $P$ is not $H_{2}$-surjective, we give estimates on the values of $g_{X^{\prime}}$ for all classes. As in Subsection 6.2, let $K^{R}_{P}$ be the kernel of the homomorphism $H_{2}(P)\to H_{2}(R)$ induced by the inclusion, and let $\iota_{P}:P\hookrightarrow X^{\prime}$ be the inclusion. Due to Lemma 6.11, we have $H_{2}(X^{\prime})=\iota_{*}(H_{2}(X))+{\iota_{P}}_{*}(K^{R}_{P})$. We prove the following estimates for $g_{X^{\prime}}$.
Theorem 7.7.
Every classes $\alpha\in H_{2}(X)$ and $\beta\in K^{R}_{P}$ satisfy
@@
The right inequality is obvious, so we prove the left inequality. Let $j:X\hookrightarrow X^{\prime\prime}$ and $j^{\prime}:X^{\prime}\hookrightarrow X^{\prime\prime}$ be the inclusions. By the definition of $K^{R}_{P}$, we see $j^{\prime}_{*}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))=j_{*}(\alpha)$, showing $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X^{\prime\prime}}(j_{*}(\alpha))$. Since $j_{*}$ preserves the genus functions due to Theorem 7.1, we obtain $g_{X^{\prime}}(\iota_{*}(\alpha)+{\iota_{P}}_{*}(\beta))\geq g_{X}(\alpha)$. ∎
As seen from Lemma 6.7, there are many examples of quasi-invertible cobordisms $P$ such that $g_{P}$ is the zero-map, and for such examples, Theorem 7.7 shows that $g_{X^{\prime}}$ is completely determined by $g_{X}$ even when $P$ is not $H_{2}$-surjective.
Under an additional assumption, the inclusion furthermore preserves the torsion-free genus functions.
@@
Suppose the homomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(X^{\prime\prime})/\mathrm{Tor}$ induced by the inclusion is injective. Then, the inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions.
Proof.
Let $A$ be a class of $H_{2}(X)/\mathrm{Tor}$, and let $\alpha$ be a class of $H_{2}(X)$ that represents $A$ and satisfies $g_{X}(\alpha)=g^{*}_{X}(A)$. By Theorem 7.1, we see
$g^{*}_{X}(A)=g_{X}(\alpha)=g_{X^{\prime}}({\iota}_{*}(\alpha))\geq g^{*}_{X^{\prime}}({\iota}_{*}(A)).$
Let $\alpha^{\prime}$ be a class of $H_{2}(X^{\prime})$ that represents ${\iota}_{*}(A)\in H_{2}(X^{\prime})/\mathrm{Tor}$ and satisfies $g_{X^{\prime}}(\alpha^{\prime})=g^{*}_{X^{\prime}}({\iota}_{*}(A))$. Proposition 6.16 implies that there exist torsion classes $\tau\in H_{2}(X)$ and $\beta\in K^{R}_{P}$ satisfying $\alpha^{\prime}={\iota}_{*}(\alpha)+{\iota}_{*}(\tau)+{\iota_{P}}_{*}(\beta)$. Theorem 7.7 and the above inequality thus imply
$g^{*}_{X^{\prime}}({\iota}_{*}(A))=g_{X^{\prime}}(\alpha^{\prime})\geq g_{X}(\alpha+\tau)\geq g^{*}_{X}(A)\geq g^{*}_{X^{\prime}}({\iota}_{*}(A)),$
showing $g^{*}_{X^{\prime}}({\iota}_{*}(A))=g^{*}_{X}(A)$. Therefore, the inclusion $\iota:X\hookrightarrow X^{\prime}$ preserves the torsion-free genus functions. ∎
We remark that the assumption of the above corollary holds if $P$ is strongly quasi-invertible. For other sufficient conditions, see Proposition 6.17.
@@ 7.2. Stabilities for algebraic inequivalences of genus functions
Throughout this subsection, we use the following setting. For $i=1,2$, let $M_{i},N_{i}$ be closed 3-manifolds, let $P_{i}$ be a quasi-invertible cobordism from $M_{i}$ to $N_{i}$, and let $Q_{i}$ be its quasi-inverse cobordism. Also, let $X_{i}$ be a connected 4-manifold having a boundary component orientedly diffeomorphic to $M_{i}$, and we denote a glued 4-manifold $X_{i}\cup_{M_{i}}P_{i}$ by $X^{\prime}_{i}$. We attach $Q_{i}$ to $X_{i}^{\prime}$ so that the submanifold $R_{i}=P_{i}\cup_{N_{i}}Q_{i}$ is a quasi-product cobordism, and we denote the resulting 4-manifold $X_{i}\cup_{M_{i}}P_{i}\cup_{N_{i}}Q_{i}$ by $X^{\prime\prime}$. Let $\iota_{i}:X_{i}\hookrightarrow X^{\prime}_{i}$ and $j_{i}:X_{i}\hookrightarrow X^{\prime\prime}_{i}$ be the inclusions. Furthermore, we define the subgroup ${{\iota_{i}}_{P_{i}}}_{*}(K^{R_{i}}_{P_{i}})$ of $H_{2}(X_{i}^{\prime})$ as in Subsection 6.2.
We will show that algebraic inequivalences of (torsion-free) genus functions are stable under attaching a quasi-invertible cobordism. Our results thus show that exotic 4-manifolds with algebraically inequivalent genus functions always yield many more (orientedly) exotic 4-manifolds by attaching quasi-invertible cobordisms.
We first prove the following theorem, which shows stabilities by taking the contrapositions.
@@
$(1)$ Suppose each ${\iota_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime}_{i})$ is an isomorphism. Then, the genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
$(2)$ Suppose that each ${\iota_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime}_{i})/\mathrm{Tor}$ is an isomorphism and that each ${j_{i}}_{*}:H_{2}(X_{i})/\mathrm{Tor}\to H_{2}(X^{\prime\prime}_{i})/\mathrm{Tor}$ is injective. Then, the torsion-free genus functions of $X_{1}^{\prime}$ and $X_{2}^{\prime}$ are algebraically equivalent if and only if those of $X_{1}$ and $X_{2}$ are algebraically equivalent.
@@
We remark that the assumptions of this theorem hold in the case where each $P_{i}$ is strongly quasi-invertible and $H_{2}$-surjective. See Propositions 6.14 and 6.17 and Lemma 6.11 for other sufficient conditions.
For general quasi-invertible cobordisms, we give a stability of algebraic inequivalences under additional assumptions. See Propositions 6.13 and 6.14 and Lemma 6.15 for sufficient conditions that the assumption below holds.
@@
Suppose the intersection form of each $X_{i}$ is non-degenerate. Suppose further that the groups $H_{2}(X_{1})$ and $H_{2}(X_{2})$ or the groups ${\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(K^{R_{2}}_{P_{2}})$ are torsion-free and that each ${j_{i}}_{*}:H_{2}(X_{i})\to H_{2}(X^{\prime\prime}_{i})$ is injective. If the genus functions of $X^{\prime}_{1}$} and $X^{\prime}_{2}$ are algebraically equivalent, then those of $X_{1}$} and $X_{2}$ are algebraically equivalent. If furthermore $H_{2}(X_{1})$ and $H_{2}(X_{2})$ are torsion-free, then the restrictions of the genus functions of $X^{\prime}_{1}$} and $X^{\prime}_{2}$ respectively to ${\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})$ and ${\iota_{P_{2}}}_{*}(K^{R_{2}}_{P_{2}})$ are also algebraically equivalent.
@@
Due to Propositions 5.3 and 5.2, the claims follow from Lemma 6.11, Proposition 6.13 and Theorems 7.1 and $7.$ ∎
For torsion-free genus functions, we give a similar stability under a much simpler assumption.
@@
Suppose the intersection form of each $X_{i}$ is non-degenerate. If the torsion-free genus functions of $X^{\prime}_{1}$} and $X^{\prime}_{2}$ are algebraically equivalent, then those of $X_{1}$} and $X_{2}$ are algebraically equivalent, and those of $X^{\prime}_{1}$} and $X^{\prime}_{2}$} restricted respectively to ${\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})/\mathrm{Tor}$ and ${\iota_{P_{2}}}_{*}(K^{R_{2}}_{P_{2}})/\mathrm{Tor}$ are algebraically equivalent.
@@
Due to Proposition 6.16, Theorem 7.7 and Corollary 7.8, we easily see that every $\alpha\in{\iota_{i}}_{*}(H_{2}(X_{i}))/\mathrm{Tor}$ and $\beta\in{\iota_{P_{1}}}_{*}(K^{R_{1}}_{P_{1}})/\mathrm{Tor}$ satisfy $g^{*}_{X^{\prime}_{i}}(\alpha)\leq g^{*}_{X^{\prime}_{i}}(\alpha+\beta)$ for each $i$ by an argument similar to the proof of Corollary 7.8. The claim thus follows from Propositions 5.3, 5.2 and 6.16 and Corollary 7.8. ∎
We remark that Theorems 7.10 and 7.11 do not hold if the non-degeneracy assumption is removed as seen from Remark 8.11.